Anle Corporation has a current stock price of $22.02 and is expected to pay a dividend of $1.20 in one year. Its expected stock price right after paying that dividend is $23.95.

a. What is​ Anle's equity cost of​ capital?

b. How much of​ Anle's equity cost of capital is expected to be satisfied by dividend yield and how much by capital​ gain?
Answer #1

a.equity cost of capital=(End value-Beginning value+Dividend)/Beginning value

=(23.95-22.02+1.20)/22.02

=14.21%(Approx).

b.Dividend yield=Dividend/Beginning value

=1.20/22.02

=5.45%(Approx).

Capital gains=(End value-Beginning value)/Beginning value

=(23.95-22.02)/22.02

=8.76%(Approx).